An object of mass m is thrown vertically upward from the surface of Mercury, where the acceleration due to gravity is gh. The initial speed is v, and the object reaches a maximum height of yr. If a similar object is thrown upward with initial speed v2 7v from the surface of Saturns moon Tethys, its maximum height is y2. The acceleration due to gravity on Tethys is 92, Express the ratio of y2 to y in terms of gl and g2- 49V181 2g y2 y1
Answer #1

using equation

v^2 = u^2 + 2gh

for surface of mercury

0 = v1^2 - 2*g1*y1

y1 = v1^2/(2*g1) ..... eq1

for surface of tethys

0 = v2^2 - 2*g2*y2

y2 = v2^2/(2*g2) ...... eq2

eq2 / eq1

y2/y1 = (49*v1^2/(2*g2) )/(v1^2/(2*g1) )

y2/y1 = (49*g1/g2)